# Kolmogorov-Arnold Fourier Networks

### v1

@@ 1 Introduction
  • By leveraging the associative property of matrix multiplication, we merge the two large matrices ($WA$ and $WB$) of KAN, thereby reducing the parameter complexity from $O(d_{\text{in}}\times d_{\text{out}}\times(G+K+3))$ to $O(d_{\text{in}}\times d_{\text{out}})$, while preserving the expressive power of the model (Arora et al., 2018; Chen et al., 2020). This approach not only effectively reduces the number of parameters but also enhances the model’s scalability in high-dimensional tasks.
@@ 3.2 Kolmogorov-Arnold Network (KAN)
where $\Phi$ is a matrix of basis functions. This formula aligns with the form of the Kolmogorov-Arnold theorem. However, in practical applications, they chose B-spline (Bach, 2016) basis functions as the basis functions $\phi_{q,p}$, and added an external activation function $SILU$ to guide the update of the KAN layer(Ramachandran et al., 2017; Elfwing et al., 2017). The formula can be expressed as
@@ 3.3 KAF: Kolmogorov-Arnold Fourier Network
where: $\tilde{\mathbf{x}}^{(l)}=\text{LayerNorm}(\mathbf{x}^{(l)})$ is the normalized input at layer $l$; $\tilde{\phi}(\cdot)$ represents the nonlinear mapping based on Random Fourier Features (RFF) (detailed in Section 3.3.2); $\mathbf{a}^{(l)},\mathbf{b}^{(l)}\in\mathbb{R}^{n}$ are learnable scaling parameters, used to modulate the contributions of GELU activation and RFF features, respectively; $\mathbf{W}^{(l)}\in\mathbb{R}^{m\times n}$ is the linear transformation weight, and $\mathbf{c}^{(l)}\in\mathbb{R}^{m}$ is the bias term.
@@
[…]
@@ A.1 KAN with B-splines: Parameter Counting
Let the B-spline(De Boor, 1972) order be $K$, and divide the domain into $G$ segments. Then: - Each segment needs $K+1$ control points, total $(G+K+1)$. - Boundary smoothness of order $K-1$ adds $2(K-1)$ virtual points. - Total per univariate spline: $G+3K-1$. (Sometimes simplified to $G+K+3$.)
@@ A.3 MLP FLOPs Computation
Standard MLP. - Linear layer $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$: $2\,d_{\mathrm{in}}\,d_{\mathrm{out}}$ FLOPs (multiply+add). - Activation: - ReLU: $1$ FLOP per output (comparison), total $d_{\mathrm{out}}$. - GELU: about $5\,d_{\mathrm{out}}$ FLOPs.
@@ B.1.1 Bochner’s Theorem and the Fourier Duality of Kernel Functions
According to Bochner’s(Rahimi & Recht, 2007a; Gradshteyn & Ryzhik, 2014) theorem , any translation-invariant positive definite kernel function $k(x,y)=k(x-y)$ can be expressed as the Fourier transform of a Gaussian measure:
[…]
@@ B.1.2 Expectation of Inner Product of Random Fourier Features
[…]
@@ B.2.2 Implementation in Backpropagation
In automatic differentiation frameworks(Baydin et al., 2018) (e.g., PyTorch), the gradient computation for RFF follows these steps: 1. Forward pass: Compute $\cos(W^{\top}x+b)$ and $\sin(W^{\top}x+b)$. 2. Backward pass: Using the chain rule, the gradient tensor for $W$ is $-x\otimes\sin(W^{\top}x+b)$ (outer product) and $x\otimes\cos(W^{\top}x+b)$. The gradient for $b$ is directly $-\sin(W^{\top}x+b)$ and $\cos(W^{\top}x+b)$. 3. Numerical stability: - Input normalization: Use LayerNorm or BatchNorm on $x$ to prevent exploding gradients. - Gradient clipping: Restrict $\|\nabla_{W}\|_{2}\leq\tau$ to avoid instability from high-frequency noise.
@@ C.4.1 Objective Function Definition
[…]
@@ C.4.2 Optimization Solution
[…]
@@
[…]
@@
[…]
@@ C.4.3 Numerical Integration Results
[…]
@@ F.1 Ablation on Cifar10
[…]
@@ F.2 Fitting experiment of sin(x) and cos(x)
To evaluate the model’s capability in approximating periodic functions, we conduct a fitting experiment on $\sin(x)$ and $\cos(x)$. Specifically, we train the model to learn the mapping $x\mapsto\sin(x)$ and $x\mapsto\cos(x)$ using a dataset of uniformly sampled points from the interval $[-20,20]$. The training objective minimizes the mean squared error (MSE) between the predicted and true values.

### v2

@@ 1 Introduction
  • By leveraging the associative property of matrix multiplication, we merge the two large matrices (WAWA and WBWB) of KAN, thereby reducing the parameter complexity from $O(d_{\text{in}}\times d_{\text{out}}\times(G+K+3))$ to $O(d_{\text{in}}\times d_{\text{out}})$, while preserving the expressive power of the model (Arora et al., 2018; Chen et al., 2020). This approach not only effectively reduces the number of parameters but also enhances the model’s scalability in high-dimensional tasks.
@@ 3.2 Kolmogorov-Arnold Network (KAN)
where $\Phi$ is a matrix of basis functions. This formula aligns with the form of the Kolmogorov-Arnold theorem. However, in practical applications, they chose B-spline (Bach, 2016) basis functions as the basis functions $\phi_{q,p}$, and added an external activation function SILUSILU to guide the update of the KAN layer(Ramachandran et al., 2017; Elfwing et al., 2017). The formula can be expressed as
@@ 3.3 KAF: Kolmogorov-Arnold Fourier Network
[…]
@@
where $W\in\mathbb{R}^{d\times m}$ and $b\in\mathbb{R}^{m}$. Here, $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product, and $\oplus$ represents the vector concatenation operation. The frequency matrix $W=[w_{1},\dots,w_{m}]$ is initialized according to an input-dimension-adaptive spectral distribution: $w_{ij}\sim\mathcal{N}(0,\sigma^{2}/d)$, where $\sigma^{2}$ represents the empirical variance of the input data. The phase shift $b$ is sampled from a uniform distribution $b_{i}\sim\mathcal{U}[0,2\pi]$, which ensures phase diversity, a crucial property for capturing local features of signals.See Appendix  B for more information on RFF convergence, gradient calculation, and initialization strategies.
@@ A.1 KAN with B-splines: Parameter Counting
Let the B-spline(De Boor, 1972) order be $K$, and divide the domain into $G$ segments. Then: - Each segment needs $K+1$ control points, total $(G+K+1)$. - Boundary smoothness of order K1K-1 adds 2(K1)2(K-1) virtual points. - Total per univariate spline: G+3K1G+3K-1. (Sometimes simplified to $G+K+3$.)
@@ A.3 MLP FLOPs Computation
Standard MLP. - Linear layer $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$: $2\,d_{\mathrm{in}}\,d_{\mathrm{out}}$ FLOPs (multiply+add). - Activation: - ReLU: 11 FLOP per output (comparison), total $d_{\mathrm{out}}$. - GELU: about $5\,d_{\mathrm{out}}$ FLOPs.
@@ B.1.1 Bochner’s Theorem and the Fourier Duality of Kernel Functions
According to Bochner’s(Rahimi & Recht, 2007a; Gradshteyn & Ryzhik, 2014) theorem , any translation-invariant positive definite kernel function k(x,y)=k(xy)k(x,y)=k(x-y) can be expressed as the Fourier transform of a Gaussian measure:
$k(x-y)=\int_{\mathbb{R}^{d}}e^{i\omega^{\top}(x-y)}p(\omega)d\omega$ (17)
@@ B.1.2 Expectation of Inner Product of Random Fourier Features
$=2\cdot\text{Re}\left(k(x-y)\right)=2k(x-y)\quad(\text{since }k(x-y)\text{ is a real-valued symmetric function}).$
@@ B.2.2 Implementation in Backpropagation
In automatic differentiation frameworks(Baydin et al., 2018) (e.g., PyTorch), the gradient computation for RFF follows these steps: 1. Forward pass: Compute $\cos(W^{\top}x+b)$ and $\sin(W^{\top}x+b)$. 2. Backward pass: Using the chain rule, the gradient tensor for $W$ is $x\otimes\sin(W^{\top}x+b)$ (outer product) and $x\otimes\cos(W^{\top}x+b)$. The gradient for $b$ is directly $\sin(W^{\top}x+b)$ and $\cos(W^{\top}x+b)$. 3. Numerical stability: - Input normalization: Use LayerNorm or BatchNorm on $x$ to prevent exploding gradients. - Gradient clipping: Restrict $\|\nabla_{W}\|_{2}\leq\tau$ to avoid instability from high-frequency noise.
@@ C.4.1 Objective Function Definition
$\mathcal{L}(\alpha)=\int_{-\infty}^{\infty}\left|S_{\text{target}}(\omega)-\alpha^{2}S_{\text{GELU}}(\omega)\right|^{2}d\omega.$ (35)
@@ C.4.2 Optimization Solution
$\mathcal{L}(\alpha)=\int_{-\infty}^{\infty}\left(1-\alpha^{2}S_{\text{GELU}}(\omega)\right)^{2}d\omega.$ (36)
@@
$\frac{d\mathcal{L}}{d\alpha}=-4\alpha\int_{-\infty}^{\infty}S_{\text{GELU}}(\omega)\left(1-\alpha^{2}S_{\text{GELU}}(\omega)\right)d\omega=0.$ (37)
@@
$\alpha_{\text{opt}}=\sqrt{\frac{\int_{-\infty}^{\infty}S_{\text{GELU}}(\omega)d\omega}{\int_{-\infty}^{\infty}S_{\text{GELU}}^{2}(\omega)d\omega}}.$ (38)
@@ C.4.3 Numerical Integration Results
$\int_{-\infty}^{\infty}S_{\text{GELU}}(\omega)d\omega\approx 0.168,\quad\int_{-\infty}^{\infty}S_{\text{GELU}}^{2}(\omega)d\omega\approx 0.062.$ (39)
@@ F.1 Ablation on Cifar10
    Effect of different num_grids values: We report the highest test accuracy for different selections of $\text{num\_grids}=9$.
@@ F.2 Fitting experiment of sin(x) and cos(x)
To evaluate the model’s capability in approximating periodic functions, we conduct a fitting experiment on $\sin(x)$ and $\cos(x)$. Specifically, we train the model to learn the mapping $x\mapsto\sin(x)$ and $x\mapsto\cos(x)$ using a dataset of uniformly sampled points from the interval [20,20][-20,20]. The training objective minimizes the mean squared error (MSE) between the predicted and true values.